OpenStudy (anonymous):

Simplify the expression 6x(5x^2 − x).

mathslover (mathslover):

\[\large{a(b-c)=ab-ac}\] So \(\large{6x(5x^2-x)=6x(5x^2)-6x(x)}\) can u do it now?

OpenStudy (anonymous):

unfortunatly no whats the next step??

mathslover (mathslover):

can u tell me what is : \[\large{6x(5x^2)=6x\times 5x^2}\]

OpenStudy (anonymous):

30x^2=30x^2?? i dont know what im doing

mathslover (mathslover):

can u tell me what is x * x^2 ?

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

0r 1x^2

mathslover (mathslover):

right it is x^3 .. hence the answer is 6 * 5 * x * x^2 = 30 x^3